Find all zeros of the polynomial. (Enter your answers as a comma-separated list. Enter all answers including repetitions.)
P(x) = x3 + 4x2 + 9x + 36


Answer by ikleyn(53751)

.

Apply grouping


      x3 + 4x2 + 9x + 36 = (x^3 + 4x^2) + (9x + 36) = x^2(x+4) + 9*(x+4) =

    = (x^2 + 9)*(x+4).


The zeroes are  3i, - 3i  (complex zeroes)  and  -4 (real zero).